OpenStudy (anonymous):

how do you find the factor of the polynomial x squared minus x minus 20

OpenStudy (anonymous):

The factorization scheme that jumped to my mind when i looked at your polynomial was ( x - a )( x + b ) = x^2 - ax + bx - ab Knowing that in your case we would need values for a and b such that b - a = -1 and ab = 20. Since 20 = 4 * 5 and 4 - 5 = -1 we have a solution at ( x - 5 )( x + 4 ) = x^2 - x - 20 as required. Does that help?